Dick is 5 years older than Jane. In 10 years twice Dick's age decrease by Jane's will be 35. Find their ages now.
Answer by tutorcecilia(2144)

Age problem:
Dick's age = Jane + 5
Jane's age = J
Twice Dick's age = 2(J + 5)
Decreased by Jane's age = -J
In 10 years = add an addition 10 years
So, putting it all together
[2(J+5) - J] + 10 years = 35 [simplify by combining like-terms]
2J+10-J+10=35
J+20=35 [solve for the J=term]
J+20-20=35-20
J=15
.
Plugging (J=15) back into Dick's and Janes age:
Dick = J + 5 = 15+5= 20 years old
Jane = J=15 years old
.
Check by plugging all of the values back into the original equation and solve:
[2(J+5) - J] + 10 years = 35

[2(15+5) - 15] + 10 years = 35
[2(20) - 15] + 10 years = 35
40-15+10=35
35=35 [checks out